OpenStudy (anonymous):

Which of the following is the slope between the two points (1, 5) and (6, 4)?

OpenStudy (ranga):

slope is difference in y values / difference in x values

OpenStudy (nurali):

(1, 5) and (6, 4) (x1,y1) and (x2,y2) m=slope Formula: m = (y2-y1)/(x2-x1) m= (4-5)/(6-1) m= -1/5